# DiffFuSR: Super-Resolution of all Sentinel-2 Multispectral Bands using Diffusion Models

### v1

@@ IV Experiments and Results
We will now evaluate the proposed pipeline (DiffFuSR) designed to super-resolve all 12 Sentinel-2 spectral bands to a 2.5-meter GSD. We first describe the datasets utilized for training the diffusion-based SR model and the learned fusion network, followed by their respective training configurations. The performance of DiffFuSR is then benchmarked against established SR and pansharpening baselines. Evaluation is conducted through a comprehensive set of quantitative metrics assessing reflectance accuracy, spectral fidelity, spatial detail, and artifact generation, complemented by qualitative visual comparisons to demonstrate the practical efficacy of the super-resolved products.
@@ IV-A1 SEN2NAIP Dataset
The SEN2NAIP dataset [3] consists of two components:
@@ IV-A2 WorldStrat Dataset
The WorldStrat dataset [13] contains HR RGB imagery acquired from the SPOT satellite. The data is globally distributed and tile-based, covering diverse geographic regions and land cover types. It is designed to support machine learning tasks in remote sensing, such as SR and domain generalization.
@@ IV-A3 Sentinel-2 Data
The Sentinel-2 data used in this work consists of Level-2A products from the Sentinel-2 mission under the Copernicus program. These products provide 12 spectral bands at spatial resolutions of 10 m, 20 m, and 60 m. All imagery has been curated to be cloud-free and atmospherically corrected, with spatial and temporal alignment across samples. This dataset serves as the LR input for fusion and SR tasks.
IV-B Training Configurations
@@
The model architecture follows the design of [14] and comprises:
  • A degradation encoder for contrastive representation learning.
  • A spatial encoder-decoder that is implemented using an RRDB-Net encoder with 23 residual blocks and 64 feature channels.
  • A diffusion UNet with 5 residual blocks and channel multipliers {1, 2, 2, 4, 8}.
  • Diffusion configuration: we use 1000 diffusion steps, with a cosine noise schedule following [14]. The model predicts the original clean image $\mathbf{x}_{0}$ directly.
  • Exponential moving average (EMA) of weights with a decay rate of 0.999.
@@
All models were trained for 1000 epochs using the Adam optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$) with an initial learning rate of $1\times 10^{-4}$ , decayed by a factor of 0.5 at epochs 200, 400, 600, and 800. We used a batch size of 8 and image patch size of 256$\times$256. Data augmentation included random cropping and flipping.
@@
Training used samples from the synthetic set of SEN2NAIP [3]. In the harmonized variant, on-the-fly histogram harmonization was applied during data loading. For unharmonized training, this step was omitted. Both configurations used four-channel multispectral inputs. Degradation modeling applied anisotropic Gaussian blur (kernel size 21, $\sigma\in[2.0,4.0]$) and additive Gaussian noise with standard deviation 25.
@@
A separate model was trained on globally distributed tiles from the WorldStrat dataset [13]. As with NAIP, four-channel input was used and the same model architecture and optimizer configuration were applied. This setup facilitates assessment of generalization across diverse geographies and domains.
IV-C Fusion Model Training with Sentinel-2
We train a cross-sensor fusion model to enhance spatial resolution in Sentinel-2 multispectral imagery by leveraging HR inputs and multi-resolution contextual information. The fusion model operates over three spatial scales corresponding to Sentinel-2 band groups: 10 m, 20 m, and 60 m resolutions.
Architecture
The model is a multi-branch variant of RRDB-Net [26], consisting of three dedicated fusion modules:
  • A 10 m branch that fuses RGB and near infrared bands (bands 2, 3 4 and 8).
  • A 20 m branch that fuses the red edge and short wave infrared bands (bands 5, 6, 7, 8A, 11, and 12).
  • A 60 m branch that fuses coastal aerosol and water vapor bands (bands 9 and 10).
Each branch uses an RRDB-Net with 5 residual blocks, 64 base channels, and group convolutions of 32. The fusion signal (i.e., the super-resolved S2 RGB bands) is concatenated with interpolated LR multispectral inputs and used to reconstruct HR outputs at each scale.
The model explicitly accounts for scale differences using a combination of bilinear interpolation, Gaussian blur (pre-smoothing) and boxcar averaging (simulating physical sensor downsampling). These methods are applied dynamically during the forward pass to simulate sensor-specific degradation and harmonize inputs across scales.
@@ IV-C Fusion Model Training with Sentinel-2
Three models are trained using separate optimizers for each spatial branch (10 m, 20 m, 60 m), each using Adam with a learning rate of $1\times 10^{-4}$. A standard L1 loss is applied independently at each scale. Training is conducted in manual optimization mode in PyTorch Lightning, allowing per-branch updates and metric tracking.
@@
During validation, the model’s outputs are compared to ground truth using the ERGAS metric, computed at each resolution. Additionally, a Gram-Schmidt pan-sharpening baseline is computed for all outputs.
@@ IV-D Baseline Methods
Existing state-of-the-art methods do not directly address the task of super-resolving all Sentinel-2 bands to a unified resolution. Since DiffFuSR is modular, we evaluate its components in two stages. First, we compare its RGB SR module against established RGB-specific SR baselines. Second, we evaluate the fusion module separately by comparing it to a suitable multispectral fusion baseline that combines super-resolved RGB with the remaining bands. This decomposition allows for a systematic assessment of each component relative to current state-of-the-art approaches.
@@ IV-D1 Super resolution baselines
We evaluate our proposed SR approach against a diverse set of existing models, representing a spectrum of training philosophies. The OpenSR baseline [3] serves as a strong reference point, particularly due to its use of degradation-based training on NAIP data. The LDM_Baseline [18] employs a latent diffusion model, applying generative priors to RGB upsampling, but suffers from instability and hallucination in complex scenes. Superimage adopts a pixel-wise regression framework optimized for spectral fidelity. Satlas [16], trained using a generative adversarial network [26] on cross-domain imagery, tends to produce sharper images but depicts poor spectral integrity. Additionally, we compare with SR4RS [17], a regression-based model targeting RGB SR via auxiliary feature learning, but which introduces spatial inconsistencies in challenging regions.
IV-D2 Fusion model baseline: Gram-Schmidt pansharpening
To benchmark the effectiveness of the learned fusion network, we implement a classical Gram-Schmidt (GS) pansharpening baseline [28, 9]. This method fuses LR multispectral bands with a HR RGB image, simulating a panchromatic band by averaging the RGB channels. The GS transformation aligns the spectral content of the MS input with the spatial detail of the pan proxy, producing a high-resolution output without learning.
For fair comparison, we apply GS using the same degraded multispectral bands and super-resolved RGB image used by our model at inference time. The GS baseline thus serves as a non-parametric alternative guided by the same spatial prior, highlighting the value added by learned fusion.
@@ IV-E1 Super Resolution
Synthesis captures how much meaningful high-frequency detail the model introduces beyond simple interpolation. Higher synthesis values indicate better recovery of fine structures such as edges, textures, and small objects that are absent in the low-resolution input.
Correctness analyzes whether the generated details are realistic. It includes: (i) hallucinations, quantifying spurious details not present in the true HR reference; (ii) omissions, measuring genuine high-resolution details that the model failed to recover; and (iii) improvements, reflecting correctly added fine-scale structures that were not visible in the low-resolution input. Ideally, hallucination and omission scores should be low, while improvement should be high.
The distance metric chosen for calculation can be either normalized difference distance, LPIPS [35], or CLIP distance [36]. Using all three gives us physical, perceptual, and semantic validation, respectively, which makes OpenSR-test uniquely strong compared to typical SR benchmarks. For full mathematical definitions and implementation details, we refer to the OpenSR-test paper [15].
@@ IV-E2 Fusion Pipeline
To evaluate the quality of the SR fusion outputs in the absence of ground truth (GT) HR Sentinel-2 imagery, we adopt a no-reference validation strategy. Specifically, we downsample the SR outputs back to the original Sentinel-2 resolution and perform comparisons against the native Sentinel-2 inputs. This indirect supervision enables us to assess fidelity and consistency across multiple quality metrics.
We conduct a quantitative comparison of different fusion methods, neural network (NN) versus GS-based fusion across the 12 Sentinel-2 bands. Evaluations are carried out using the following metrics:
  • R2: Coefficient of determination, indicating how well predicted reflectance values approximate the actual values. Higher R2-values is better, indicating a strong alignment between the predicted and actual values.
  • Cross-Correlation: Measures the linear relationship between the predicted and original signal. Values close to one indicate strong similarity.
  • SSIM (Structural Similarity Index): Assesses perceptual similarity in terms of structure and luminance. Higher values denote more faithful reconstructions [37].
  • PSNR (Peak Signal-to-Noise Ratio): Evaluates the ratio between the maximum possible signal and the power of corrupting noise. Higher PSNR suggests better fidelity [37].
  • MSE (Mean Squared Error): Captures the average squared difference between the downsampled SR image and the original Sentinel-2 image. Lower values are preferred.
  • ERGAS (Relative Dimensionless Global Error in Synthesis): A global error metric specific to remote sensing fusion tasks. Lower values indicate better spectral and spatial consistency [19].
@@ IV-F1 Super resolution
Figs. 4 and 5 illustrate qualitative comparisons of the proposed diffusion-based SR model trained on three different datasets: harmonized NAIP, unharmonized NAIP, and WorldStrat. Each figure is organized into three major rows (one per dataset), where the top row shows reconstructed images, including the LR input, downsampled version, SR output, harmonized SR output, and the HR reference. The second row visualizes per-pixel evaluation maps for reflectance consistency, spectral consistency (SAD), omission space, hallucination space, and improvement space, offering visual insight into the model’s behavior.
@@
In Fig. 4, which features a forested scene with green tree cover and bare land, the harmonized NAIP model clearly yields the sharpest and most visually realistic reconstruction. Both the SR image and the harmonized version closely resemble the HR reference, with fine details preserved. This is supported by the quantitative overlays, where reflectance and spectral consistency metrics are lowest (L1: 0.0021, SAD: 1.3884) and omission and hallucination errors are moderate. The unharmonized NAIP variant performs reasonably well but introduces more spectral error and slightly higher hallucination (SAD: 1.5671, HA: 0.4705). The WorldStrat-trained model appears perceptually blurrier and less structurally aligned, despite having similar reflectance L1, and its SAD score (2.10) and omission space distance (0.5530) indicate weaker spectral fidelity. This case highlights the strength of NAIP-derived training, especially in harmonizing the reconstruction of vegetated regions. The likely explanation is that there is a heavy focus on areas with vegetation in NAIP dataset, whereas the WorldStrat dataset has a slight urban area tilt.
Conversely, Fig. 5 presents an urban scene with buildings, where the performance trend shifts. Here, the WorldStrat-trained model produces the most visually convincing output with edges being sharp, and structural details like roof boundaries and roads being clearly resolved. Despite having a slightly higher spectral error (SAD: 1.8039), its visual quality surpasses that of both NAIP-trained models. The harmonized NAIP model underperforms in this case, with noticeable blurring and misalignment around fine structures, and a higher hallucination score (0.4900). This result demonstrates that while NAIP performs well in vegetated regions, it may struggle in built-up areas where WorldStrat’s diverse urban coverage offers an advantage.
@@ IV-F2 Fusion module
Fig. 6 evaluates the proposed learnable fusion model. Here, we only use the harmonised NAIP-based SR model to create the RGB super-resolved image that is further used for fusion. The columns represent the LR input, GS fusion, and our neural network-based fusion, applied to 60 m bands and visualized in natural and false color. Compared to GS, which often distorts the structure and exaggerates spectral contrast, the fusion model output is visually more consistent with the low-resolution signal while enhancing spatial detail. Particularly in the 60 m bands, the fusion model preserves smooth gradients and subtle features that are entirely lost in GS. These results confirm that the learned fusion pipeline generalizes well and substantially outperforms classical pansharpening.
Together, these examples highlight the importance of dataset diversity and the advantage of learning-based fusion. NAIP excels in natural environments, WorldStrat performs better in urban settings, and neural fusion consistently surpasses traditional methods.
Low Resolution Gram Schmidt Fusion Network
Figure 6: Fusion output comparisons across 60 m atmospheric bands (top row), natural color (middle), and false color composites (bottom). Each column shows the LR input, GS pansharpening baseline, and our learned neural fusion output. The neural model better preserves semantic structure and avoids the spectral distortions typical in GS fusion, especially in low-texture regions.
@@ IV-G Quantitative Evaluation
To assess the efficacy of the proposed SR pipeline, we conducted a comprehensive quantitative evaluation across three distance metrics: Normalized Difference, CLIP, and LPIPS [15]. These metrics collectively assess reflectance preservation, spectral and spatial fidelity, and perceptual realism. In addition, we summarize the performance of each method using average rank. To obtain the average rank, each model is ranked individually for every metric: for metrics where lower values indicate better performance (Reflectance, Spectral, Spatial, Hallucination, and Omission), ranks are assigned from lowest (rank 1) to highest (rank 8); for metrics where higher values are better (Synthesis and Improvement), ranks are assigned from highest (rank 1) to lowest (rank 8). The average rank for each model is then computed as the arithmetic mean of its ranks across all seven metrics, providing a single aggregated measure of overall performance.
@@ IV-G1 Performance on Normalized Difference Metrics
The results in Table I show that the proposed diffusion model, trained on harmonized NAIP data, achieves best-in-class performance in multiple key metrics, including average rank. Specifically, it achieves the lowest reflectance error ($0.0024$), highest spectral fidelity ($1.1103$), and the second-best spatial alignment ($0.0052$). Its synthesis score ($0.0083$) and omission ($0.3543$) are high among real competitors, while hallucination errors remain moderate compared to baselines. Notably, our model’s improvement score ($0.1876$) confirms its strong ability to enrich LR inputs with high-frequency, realistic details.
@@ IV-G3 Insights from LPIPS Metric
As illustrated in Table III, LPIPS-based analysis shows that the proposed method not only achieves strong perceptual fidelity but also minimizes over-synthesis. It delivers a moderate hallucination score ($0.1149$), confirming that our model avoids generating too many spurious details. Additionally, it achieves the best improvement score ($0.5383$), reaffirming its effectiveness in reconstructing rich, high-frequency features from low-resolution imagery.
It is interesting to note that the DiffFuSR-based models achieve the best average ranking across the three Tables.
@@ IV-G4 Fusion-based enhancement of multispectral bands
To super-resolve all 12 Sentinel-2 bands, we combine RGB upsampling with learned fusion models trained to upscale 10 m, 20 m, and 60 m bands to a unified 2.5 m resolution. Table IV presents a comparison with classical GS pansharpening. Across all data sources, our neural network (NN)-based fusion dramatically outperforms GS in terms of ERGAS, PSNR, and SSIM. For instance, ERGAS drops from over $2900$ (GS) to under $95$ (NN), PSNR increases from  22 dB to  32 dB, and SSIM rises from  0.65 to  0.88.
TABLE IV: Quantitative Comparison of Super-Resolution Fusion Methods (NN vs. GS) for 12 Sentinel-2 Bands (Best Across Dataset is Bold)
Data Source Method R2 Cross-Correlation SSIM PSNR MSE ERGAS
NAIP Harmonized GS -0.343 0.905 0.648 22.76 1002.61 2972.76
NN 0.804 0.981 0.877 32.34 143.01 94.83
NAIP Unharmonized GS -0.306 0.908 0.656 22.95 977.47 3074.05
NN 0.805 0.981 0.876 32.10 142.45 95.42
WorldStrat GS -0.341 0.905 0.648 22.75 1001.48 57515.33
NN 0.806 0.981 0.878 32.27 141.96 94.68

### v2

@@ IV-A1 SEN2NAIP Dataset
The SEN2NAIP dataset [aybar2024sen2naip] consists of two components:
@@ IV-A2 WorldStrat Dataset
The WorldStrat dataset [cornebise2022open] contains HR RGB imagery acquired from the SPOT satellite. The data is globally distributed and tile-based, covering diverse geographic regions and land cover types.
@@ IV-A3 Sentinel-2 Data
The Sentinel-2 data used in this work consists of Level-2A products from the Sentinel-2 mission under the Copernicus program. These products provide 12 spectral bands at spatial resolutions of 10 m, 20 m, and 60 m. All imagery has been curated to be cloud-free and atmospherically corrected, with spatial and temporal alignment across samples.
IV-B Super Resolution Model Training
@@
The model architecture follows the design of Wu et al. [wu2023conditional] and comprises:
  • A degradation encoder for contrastive representation learning. (2.6M parameters)
  • A spatial encoder-decoder that is implemented using an RRDB-Net encoder with 23 residual blocks and 64 feature channels.(16.7M parameters)
  • A diffusion UNet with 5 residual blocks and channel multipliers {1, 2, 2, 4, 8}. (148.3M parameters)
  • Diffusion configuration: we use 1000 diffusion steps, with a cosine noise schedule [wu2023conditional]. The model predicts the original clean image $\mathbf{x}_{0}$ directly.
  • Exponential moving average of weights with a decay rate of 0.999.
@@
All models were trained for 1000 epochs using the Adam optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$) with an initial learning rate of $1\times 10^{-4}$ , decayed by a factor of 0.5 at epochs 200, 400, 600, and 800. We used a batch size of 8 and image patch size of $\times256$. Data augmentation included random cropping and flipping.
@@
Training used samples from the synthetic set of SEN2NAIP [aybar2024sen2naip]. In the harmonized variant, on-the-fly histogram harmonization was applied during data loading. For unharmonized training, this step was omitted. Both configurations used four-channel multispectral inputs. Degradation modeling applied anisotropic Gaussian blur (kernel size 21, $\sigma\in[2.0,4.0]$) and additive Gaussian noise with standard deviation 25.
@@
A separate model was trained on globally distributed tiles from the WorldStrat dataset [cornebise2022open]. As with NAIP, four-channel input was used and the same model architecture and optimizer configuration were applied.
IV-C Fusion Model Training
Model Architecture
The proposed fusion model is implemented as three dedicated MTFNetFusion modules, one per resolution group (10 m, 20 m, 60 m):
  • A 10 m branch that fuses RGB and near infrared bands (bands 2, 3 4 and 8). (202k parameters)
  • A 20 m branch that fuses the red edge and short wave infrared bands (bands 5, 6, 7, 8A, 11, and 12). (202k parameters)
  • A 60 m branch that fuses coastal aerosol and water vapor bands (bands 9 and 10). (202k parameters)
Each branch follows a MTF-GLP design, as specified in Sec. III-B.
@@ IV-C Fusion Model Training
Three models are trained using separate optimizers for each spatial branch (10 m, 20 m, 60 m), each using Adam with a learning rate of $1\times 10^{-4}$.
@@
Validation is performed using ERGAS at each resolution. Additionally, multiple pan-sharpening baselines are computed for comparison.
Inference Time
The inference time for the diffusion model is 1.74 seconds, and for the fusion model is 0.0161 seconds.
@@ IV-D Baseline Methods
Existing state-of-the-art methods do not directly address the task of super-resolving all Sentinel-2 bands to a unified resolution. Since DiffFuSR is modular, we evaluate its components in two stages. First, we compare its RGB SR module against established RGB-specific SR baselines. Second, we evaluate the fusion module separately by comparing it to a suitable multispectral fusion baseline that combines super-resolved RGB with the remaining bands.
@@ IV-D1 Super resolution baselines
The proposed SR approach is evaluated against a diverse set of existing models. The OpenSR baseline [aybar2024sen2naip] serves as a strong reference point, particularly due to its use of degradation-based training on NAIP data. The LDM_Baseline [10887321] employs a latent diffusion model, applying generative priors to RGB upsampling. Superimage adopts a pixel-wise regression framework optimized for spectral fidelity. We compare against Satlas [wolters2023zooming], a GAN-based model trained on cross-domain imagery. Additionally, we compare with SR4RS [cresson2022sr4rs], a regression-based model targeting RGB SR via auxiliary feature learning.
IV-D2 Fusion model baseline
To benchmark the effectiveness of the proposed fusion network (GLP-NN), we implement Gram-Schmidt (GS) pansharpening  [vivone2014critical], the intensity-hue-saturation (IHS) transform, principal component analysis (PCA), and the classical GLP with MTF filters and full-scale regression injection model (GLP) method  [vivone2018full]. Furthermore, following [deng2022machine], we include their state-of-the-art FusionNet as a learning-based baseline. Although FusionNet was originally designed for generic pansharpening, we retrain it under the same conditions as our proposed fusion module to ensure a fair comparison.
@@ IV-E1 Super Resolution
Synthesis captures how much meaningful high-frequency detail the model introduces beyond simple interpolation.
Correctness analyzes whether the generated details are realistic. It includes: (i) hallucinations, quantifying spurious details not present in the true HR reference; (ii) omissions, measuring genuine HR details that the model failed to recover; and (iii) improvements, reflecting correctly added fine-scale structures that were not visible in the low-resolution input.
The distance metric chosen for calculation can be either normalized difference distance, LPIPS [zhang2018unreasonable], or CLIP distance [radford2021learning]. For full mathematical definitions and implementation details, we refer to the OpenSR-test paper [opensrtest].
@@ IV-E2 Fusion Pipeline
To evaluate the quality of the SR fusion outputs in the absence of ground truth HR Sentinel-2 imagery, we adopt a no-reference validation strategy. Specifically, we downsample the SR outputs back to the original Sentinel-2 resolution and perform comparisons against the native Sentinel-2 inputs. Evaluations are carried out using the following metrics: coefficient of determination (R2), cross-correlation, Structural Similarity Index (SSIM) [blau2018perception], Peak Signal-to-Noise Ratio (PSNR) [blau2018perception], Mean Squared Error (MSE), and ERGAS (Relative Dimensionless Global Error in Synthesis) [wald2002data].
@@ IV-F1 Super resolution
Figs. 5 and 6 illustrate qualitative comparisons of the proposed SR model trained on three different datasets: harmonized NAIP, unharmonized NAIP, and WorldStrat. Each figure is organized into three major rows (one per dataset), where each row shows images, including the LR input, downsampled version, SR output, harmonized SR output, and the HR reference.
@@
In Fig. 5, which features a forested scene with green tree cover and bare land, the harmonized NAIP model clearly yields the sharpest and most visually realistic reconstruction. Both the SR image and the harmonized version closely resemble the HR reference, with fine details preserved. This is supported by the quantitative values, where reflectance and spectral consistency metrics are lowest (L1: 0.0021, SAD: 1.3884) and omission and hallucination errors are moderate. The unharmonized NAIP variant performs reasonably well but introduces more spectral error and slightly higher hallucination (SAD: 1.5671, HA: 0.4705). The WorldStrat-trained model appears perceptually blurrier and less structurally aligned, despite having similar reflectance L1, and its SAD score (2.10) and omission space distance (0.5530) indicate weaker spectral fidelity. The likely explanation is that there is a heavy focus on areas with vegetation in NAIP dataset, whereas the WorldStrat dataset has a slight urban area tilt.
Conversely, Fig. 6 presents an urban scene with buildings, where the performance trend shifts. Here, the WorldStrat-trained model produces the most visually convincing output with edges being sharp, and structural details like roof boundaries and roads being clearly resolved. Despite having a slightly higher spectral error (SAD: 1.8039), its visual quality surpasses that of both NAIP-trained models. The harmonized NAIP model underperforms in this case, with noticeable blurring and misalignment around fine structures, and a higher hallucination score (0.4900).
The SR images in Fig. 5 and Fig. 6 are histogram-matched with the HR ground truth image to obtain SRharm which is the harmonized SR output. This additional SR output is used as a diagnostic tool: a minimal visual change post-matching indicates high intrinsic spectral fidelity.
@@ IV-F2 Fusion module
Fig. 7 evaluates the proposed learnable fusion model. Here, we only use the harmonised NAIP-based SR model to create the RGB super-resolved image that is further used for fusion. The columns represent the Low Resolution input, Generalized laplacian pyramid (GLP) baseline method [vivone2018full], our GLP inspired neural network, intensity-hue-saturation (IHS) transform, and Fusion network baseline [deng2022machine] and Gram Schmidt [dalla2015global] method applied to 60 m bands and visualized in RGB, where the red and green channels use the Water Vapour band (B10) and the blue channel uses the Coastal Aerosol band (B1). Compared to the baselines, which often distort the structure and exaggerate spectral contrast, our GLP inspired fusion networks output is visually more consistent with the low-resolution signal while enhancing spatial detail.
Low Resolution GLP [vivone2018full] GLP NN (ours)
IHS Fusion Network [deng2022machine] Gram Schmidt [dalla2015global]
Figure 7: Fusion output comparisons across 60 m atmospheric bands. The figure show the Low Resolution input, Generalized laplacian pyramid (GLP) baseline method [vivone2018full], our GLP inspired neural network, intensity-hue-saturation (IHS) transform, and Fusion network baseline [deng2022machine] and Gram Schmidt [dalla2015global] method.
@@ IV-G Quantitative Evaluation
To obtain the average rank, each model is ranked individually for every metric: for metrics where lower values indicate better performance (Reflectance, Spectral, Spatial, Hallucination, and Omission), ranks are assigned from lowest (rank 1) to highest (rank 8); for metrics where higher values are better (Synthesis and Improvement), ranks are assigned from highest (rank 1) to lowest (rank 8). The average rank for each model is then computed as the arithmetic mean of its ranks across all seven metrics, providing a single aggregated measure of overall performance.
@@ IV-G1 Performance on Normalized Difference Metrics
The results in show that the proposed diffusion model, trained on harmonized NAIP data, achieves best-in-class performance in multiple key metrics, including average rank (Table I). Specifically, it achieves the lowest reflectance error ($0.0024$), highest spectral fidelity ($1.1103$), and the second-best spatial alignment ($0.0052$). Its synthesis score ($0.0083$) and omission ($0.3543$) are high among real competitors, while hallucination errors remain moderate compared to baselines.
@@ IV-G3 Insights from LPIPS Metric
LPIPS-based analysis shows that the proposed method not only achieves strong perceptual fidelity but also minimizes over-synthesis (Table III). It delivers a moderate hallucination score ($0.1149$), confirming that our model avoids generating too many spurious details. Additionally, it achieves the best improvement score ($0.5383$), reaffirming its effectiveness in reconstructing rich, high-frequency features from low-resolution imagery.
@@ IV-G4 Fusion-based enhancement of multispectral bands
To super-resolve all 12 Sentinel-2 bands, we combine RGB upsampling with learned fusion model trained to upscale 10 m, 20 m, and 60 m bands to a unified 2.5 m resolution. Table IV presents a comparison with classical baselines and the FusionNet baseline [deng2022machine]. Since HR ground truth is not available for all bands, performance must be assessed indirectly. To this end, we downsample the super-resolved outputs back to the native Sentinel-2 resolution and compare them there. Overall, our approach outperforms all classical baselines except for the GLP method [vivone2018full], which remains highly competitive, particularly on the 20 m and 60 m bands (see Table IV).
TABLE IV: Performance of the WorldStrat-trained super-resolution model coupled with our proposed GLP-NN fusion model, evaluated on bands at 10 m, 20 m, and 60 m resolution. The baselines include Generalized Laplacian Pyramid (GLP) [vivone2018full], Gram Schmidt (GS) [dalla2015global], FusionNet [deng2022machine], BOLD indicates best and UNDERLINED indicates second best
Resolution Method PSNR SSIM MSE R2 NCC ERGAS
10m (Ratio 4) GLP 27.95 0.852 12.08 0.977 0.997 2.766
GLP NN 29.57 0.904 8.56 0.984 0.998 2.260
IHS 27.90 0.870 15.82 0.972 0.996 2.583
GS 16.75 0.678 188.99 0.717 0.976 10.141
PCA 17.53 0.118 254.67 0.552 0.941 10.627
Fusion Net 29.48 0.898 8.62 0.984 0.998 2.275
20m (Ratio 8) GLP 24.39 0.836 15.51 0.952 0.998 0.789
GLP NN 22.01 0.757 26.26 0.916 0.997 1.030
IHS 17.92 0.571 84.89 0.782 0.990 1.820
GS 4.38 0.442 1584.73 -3.885 0.908 7.772
PCA 15.09 0.195 257.04 0.295 0.970 2.948
Fusion Net 21.10 0.767 32.86 0.904 0.997 1.288
60m (Ratio 24) GLP 27.92 0.957 3.19 0.997 0.999 0.188
GLP NN 23.65 0.909 5.05 0.995 0.999 0.322
IHS 22.63 0.828 21.25 0.982 0.997 0.312
GS 7.42 0.515 847.22 0.310 0.937 1.937
PCA 17.79 0.473 86.33 0.926 0.987 0.689
Fusion Net 23.16 0.918 7.21 0.992 0.999 0.380
IV-H Ablations
Our experiments show that harmonization provides a consistent, albeit modest, improvement in performance.
A more critical factor is the choice of the degradation kernel. Instead of fixing the Gaussian blur kernel to a single value (e.g., $\sigma=3.0$), we replace it with a blind kernel sampled across a range ($\sigma\in[2,4]$). This change produces a dramatic improvement: with a fixed kernel the model performance degrades sharply, while the blind kernel yields much more robust results. The importance of this choice is evident both quantitatively (Table V) and qualitatively (see Figure 8). Without the blind kernel, the entire pipeline nearly collapses, highlighting its central role in the training setup.
TABLE V: Ablation metrics across LPIPS, Normalized Difference, and CLIP. (↓ lower is better; ↑ higher is better). The selected models for comparison are trained on NAIP data using three setting i.e. Harmonized data with anisotropic blind kernel(Harm), Unharmonized data without harmonization and Harmonized data with a fixed kernel
LPIPS Normalized Difference CLIP
Metric Harm Unharm Harm (Fix Kernel) Harm Unharm Harm (Fix Kernel) Harm Unharm Harm (Fix Kernel)
Reflectance $\downarrow$ $0069\pm 0.0031 0.0$$109\pm 0.0030 0.0$$329\pm 0.0187 0.0$$024\pm 0.0009 0.0$$038\pm 0.0008 0.0$$122\pm 0.0067 0.0$$069\pm 0.0032 0.0$$109\pm 0.0030 0.0$$0329\pm 0.0187$
Spectral $\downarrow$ $1.6738\pm 0.8602$ $2.0734\pm 1.0205$ $7.3211\pm 3.1621$ $1.1103\pm 0.6197$ $1.2320\pm 0.7256$ $1.5652\pm 0.9009$ $1.6739\pm 0.8602$ $2.0734\pm 1.0205$ $7.3211\pm 3.1621$
Spatial $\downarrow$ $0060\pm 0.0097 0.0$$061\pm 0.0096 0.0$$132\pm 0.0123 0.0$$052\pm 0.0092 0.0$$053\pm 0.0093 0.0$$130\pm 0.0118 0.0$$061\pm 0.0098 0.0$$061\pm 0.0096 0.0$$0132\pm 0.0123$
Synthesis $\uparrow$ $0237\pm 0.0120 0.0$$245\pm 0.0118 0.0$$570\pm 0.0325 0.0$$083\pm 0.0043 0.0$$086\pm 0.0042 0.0$$187\pm 0.0101 0.0$$238\pm 0.0120 0.0$$245\pm 0.0118 0.0$$0570\pm 0.0325$
Hallucination $\downarrow$ $0.1149\pm 0.1020$ $0.1434\pm 0.1158$ $0.5222\pm 0.1747$ $0.4580\pm 0.1020$ $0.4927\pm 0.1095$ $0.7463\pm 0.0521$ $0.2883\pm 0.2333$ $0.3272\pm 0.2400$ $0.3810\pm 0.1223$
Omission $\downarrow$ $0.3466\pm 0.1493$ $2846\pm 0.1508 0.0$$0508\pm 0.0368$ $0.3543\pm 0.0836$ $0.3373\pm 0.0870$ $0.1376\pm 0.0373$ $0.3358\pm 0.2214$ $0.2929\pm 0.2169$ $0.1003\pm 0.0289$
Improvement $\uparrow$ $0.5383\pm 0.1240$ $0.5720\pm 0.1242$ $0.4269\pm 0.1611$ $0.1876\pm 0.0344$ $0.1701\pm 0.0351$ $0.1160\pm 0.0207$ $0.3759\pm 0.1444$ $0.3799\pm 0.1475$ $0.5186\pm 0.1202$
NAIP Harmonized
NAIP UnHarmonized
NAIP Harm, Fixed Kernel
Figure 8: Visual comparison of Ablations of our SR method on OpenSR-test  [opensrtest] test image. Among all methods, DiffFuSR with NAIP harmonized produces the sharpest reconstruction with fewer artifacts, using fixed kernel instead of blind kernel severly degrades the results.